OpenStudy (anonymous):

What is the degrees Celsius if the temperature is -15°F? Show all work. Remember to label your answer. If necessary, round your final answer to the nearest tenth.

OpenStudy (31356):

-26.1111

OpenStudy (31356):

And then round it to the nearest tenth.

OpenStudy (31356):

Medal?

OpenStudy (anonymous):

@31356 You didnt show your work

OpenStudy (loser66):

(C/5) = ((F -32)/9 C = (5/9)(F-32) C = (5/9)(-15-32) = -26.111

OpenStudy (31356):

Here's How: Take the temperature in Fahrenheit subtract 32. Divide by 1.8. The result is degrees Celsius.

OpenStudy (loser66):

medal free.!! feel free to NOT give me medal

OpenStudy (anonymous):

@Loser66 Thank you love, you get a cookie..

OpenStudy (31356):

Lol @Loser66

OpenStudy (loser66):

I practice on other's question. It's new to me, just study from Physics class. hehehe. I have to give the Asker my medal for free practice problem

OpenStudy (anonymous):

Yaaaaaay.

OpenStudy (anonymous):

@Loser66 Im putting up another practice problem, check it out if youd like.